3.2 + (x + 5.1)=
Simplify the expression.
Explain each step.
Which properties of math did you use.
My son originally got 8.3x but it was wrong. Thanks in advance.

Answer by richard1234(7193)

Correct is 8.3 + x. You are just adding three terms and since addition is commutative and associative, it does not matter what order you add. A more thorough solution:

3.2 + (x + 5.1) = 3.2 + (5.1 + x) (by commutativity)
= (3.2 + 5.1) + x (by associativity)
= 8.3 + x